Find centroid of a triangle, mid-points of whose sides are (1, 2, - 3), (2, 0, 1) and (- 1, 1, - 4).

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
let the vertices of a triangle are (x1,y1,z1) (x2,y2,z2) (x3,y3,z3)
given (x1+x2)/2=1
(x2+x3)/2=2
(x3+x1)/2=-1
by adding the above equations we get x1+x2+x3=2
simlarly y1+y2+y3=3/2
z1+z2+z3=-3
Hence centroid of triangle is (2/3 , ½ , -1)
Thanks & Regards
Shaik Aasif Ahamed
askIITians Faculty